Birch Company normally produces and sells 50,000 units of RG-6 each month. The selling price is $20 per unit, variable costs are $10 per unit, fixed manufacturing overhead costs total $175,000 per month, and fixed selling costs total $32,000 per month. Employment-contract strikes in the companies that purchase the bulk of the RG-6 units have caused Birch Company’s sales to temporarily drop to only 10,000 units per month. Birch Company estimates that the strikes will last for two months, after which time sales of RG-6 should return to normal. Due to the current low level of sales, Birch Company is thinking about closing down its own plant during the strike, which would reduce its fixed manufacturing overhead costs by $46,000 per month and its fixed selling costs by 10%. Start-up costs at the end of the shutdown period would total $15,000. Because Birch Company uses Lean Production methods, no inventories are on hand. Required: 1. What is the financial advantage (disadvantage) if Birch closes its own plant for two months? 2. Should Birch close the plant for two months? 3. At what level of unit sales for the two-month period would Birch Company be indifferent between closing the plant or keeping it open?
Answer #1

Solution 1:

Sales Price per unit = $20

Variable Cost per Unit = $ 10

Contribution Per unit = $20 - $10 = $10

Net Income if plan continues in strike period of 2 Months:

Net Income = Total Contribution - Fixed manufacturing overhead – Fixed Selling Cost

                      = (20000*10) – (175000*2) – (32000*2) = ($214,000)

Net Income if plant shut down in strike period of 2 months:

Revised fixed manufacturing cost = $175,000 - $46,000 = $129,000 per month

Revised fixed selling cost = $32,000 – ($32,000*10%) = $ 28,800 per month

Start up cost at the end of shut down = $15,000

Net Income = Total Contribution - Fixed manufacturing overhead – Fixed Selling Cost – Startup Cost

                       = 0 – (129000*2) – (28800*2) - 15000 = ($330,600)

Hence if plant is closed for 2 months than net income will decrease by $116,600 ($214,000 - $330,600).

Solution 2:

Birch should not close the plant for 2 months.

Solution 3:

If our loss will remain constant $330,600 at opening of plant and closing of plant then Birch Company will be indifferent. Let x is no of units sold at which Birch company will be indifferent between closing the plant and keeping it open.

It means

Loss at the time of opening the plant = $330,600

10x – (175000*2) – (32000*2) = -330,600

10x - 350000 – 64000 = -330600

10x = 83,400

X = 8340 units

Hence at 8340 units of sales for 2 months period Birch company be indifferent between closing the plant and keeping it open.